OpenStudy (anonymous):

The sum of two numbers is 115. Their difference is 15. What is the larger number?

OpenStudy (bahrom7893):

x+y=115 x-y=15

OpenStudy (lgbasallote):

x+ y = 115 x - y = 15 system (elimination) 2x = 130 x = 65 y = 50

OpenStudy (bahrom7893):

system eliminated. mission accomplished.

OpenStudy (anonymous):

So Which one would be the largest number?

OpenStudy (lgbasallote):

60 is larger...50 is smaller

OpenStudy (lgbasallote):

65* is largest

OpenStudy (anonymous):

Thank you very much